Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Fiori, Tringali, Paoletti, Bersanetti, Gherardini - 7:23 Thursday 11 December 2025 (68350) Print this report
Magnetic line injection

This evening, we resumed the magnetic line injection (elog #62843). Due to some issues with the Automatic Injection code, we used the ENVnoise process from the VPM interface instead.

We tuned the amplitude of the lines, starting with the WE building.

All the lines are injected simultaneously. The frequencies and amplitudes (sent to DAC) are reported in the table below. 

Frequency (Hz) Amplitude (V)
2.5  0.4
3.6  0.32
4.6 0.32
5.7  0.36
6.7  0.36
7.4 0.36
8.9  0.4
10.5 0.22
11.9 0.28
13.5 0.34
15.3  0.38
17.5 0.4
19.1 0.48
22.5 0.54
24.5  0.58 

 

During the tuning, the interferometer unlocked and the auto-relock process got stuck. We contacted Diego, who fixed the problem.

*** We left the WEB magnetic line injection running for the entire night, and tomorrow morning the injection will be stopped manually. ***

Comments to this report:
Tringali, Fiori - 7:27 Thursday 11 December 2025 (68352) Print this report

The magnetic line injections were stopped at 06:18:44 UTC.

fiori - 14:40 Thursday 11 December 2025 (68357) Print this report

A first look at data show that:

injected lines at 6.7 Hz and above are all well visible in hrec (using 1000s spectral resolution)

Yet, the choice of setting lines's frequency at multiples of 0.1 was not good, because of the presence (forgotten) of an underline 0.1Hz spaced comb which is largely coherent between hrec and magnetometers. Therefore, we would possibly repeat the injection with a better choice of frequencies. 

Injected lines below 6.7 Hz are not seen in hrec (although there is margin of improving, for example removing loud glitches...). These freqeuncies should be amplified. Indeed there is margin (before saturating the amplifier) to safely double their amplitude. 

Images attached to this comment
Tringali, Fiori, Berni - 22:55 Thursday 11 December 2025 (68362) Print this report

Due to the comb lines spaced by 0.1 Hz, we changed the frequencies of all injected lines and increased the amplitude of the first four lines in order to improve their SNR.

*** The injection started at 21:02:51 UTC in the WE building, and it will be stopped manually tomorrow morning****

Frequency (Hz) Amplitude (V)
2.577  0.8
3.677  0.64
4.677 0.64
5.777  0.64
6.777  0.36
7.477 0.36
8.977  0.4
10.577 0.22
11.977 0.28
13.577 0.34
15.377  0.38
17.577 0.4
19.177 0.48
22.577 0.54
24.577  0.58 

 

Tringali - 10:09 Friday 12 December 2025 (68365) Print this report

The WEB magnetic line injections were stopped this morning at 08:29:12 UTC.

fiori, tringali, paoletti - 17:02 Friday 12 December 2025 (68372) Print this report

We had a look at tonight injections and compared them with those of the previous night

The examined periods are:

17000 s starting from Dec 11, 01:00 

17000 s starting from Dec 11, 22:20 

We used a spectra resolution of 2000s.

The first dataset is polluted by a 0.1 Hz spaced  comb, which is present in the Voltage and current monitors of the large coil amplifier as well as in the witness magnetometer (placed close to the WE tower).  We suspect this might be related to the injected lines being multiples of 0.1 Hz.  The problem with this is that Hrec itself is polluted by a 0.1 Hz comb, although it is not related with our injected noise. Infact, as shown in Figure 16 (example around the injected line 10.5Hz) the same 0.1 Hz comb was present also when not injecting (purple). 

Therefore, in the second set of injections we moved the lins frequencies at **.*77 Hz. This time we noticed a somehow larger spectral noise "floor" in the Voltage, Current monitors and magnetometers as well, which resembles a comb spaced by 0.77 Hz. However, this turns out to be better, because no such comb is present in Hrec a priori.

Going through the injected lines one by one (Figures 1 to 15), it looks that we have achieved a good measurement of CF at most of them. Missing ones for now are 3.6, 4.6, 5.7 Hz, which might be recovered improving the analysis.

We do not have a good explanation for the extra noise  produced by the amplifier, it seems related to a non linearity in the amplifier itself, since the spectrum of the injected noise (ENV_NOISE_MAG_WEB) does not show this features.

 

Images attached to this comment
Tringali - 17:57 Sunday 11 January 2026 (68465) Print this report

The issue observed in December with the automated magnetic line injection has been resolved. The Python code for the magnetic line injections has been updated.
In particular, improvements were implemented in the .py script to enhance the parsing and handling of the configuration (.ini) file.

It was also clarified that the Amp and Freq parameter lists must always have the same length; otherwise, the code terminates without producing explicit error messages (this indicates that the code could be further improved).

A test injection was performed using the ENV Metraton node, injecting two magnetic lines at 10 Hz and 20 Hz. The injections were carried out first on WEB and then on NEB.

The injection procedure worked as expected.

Images attached to this comment
Tringali, Gherardini - 10:36 Thursday 22 January 2026 (68547) Print this report

Yesterday night (Jan 21), the magnetic line injections were performed via the ENV METRATON node in NEB and WEB.
Regarding the CEB magnetic lines injection, this morning I found that it was still ongoing. In principle, it should have been completed during the night, but there was an interferometer unlock (during NEB injection). I stopped the CEB lines injection this morning since the interferometer had been unlocked to allow the start of the scheduled shift activities.

The output files are /virgoData/NoiseInjections/MagneticInjectionsO4/output:

  • MagneticLine_NOISE_MAG_CEB-1453096947.txt
  • MagneticLine_NOISE_MAG_NEB-1453060711.txt
  • MagneticLine_NOISE_MAG_NEB-1453073344.txt
  • MagneticLine_NOISE_MAG_WEB-1453085146.txt
Images attached to this comment
Tringali, Fiori - 20:14 Saturday 24 January 2026 (68563) Print this report

Before starting the magnetic line and magnetic sweep injections today, we performed a few tests at WEB, since the last lines injection (#68547) did not work properly.

We injected a magnetic sweep via the Metatron node and a magnetic line via VPM; however, in both cases no magnetic field was generated. Both the coil switch-on and the DAC signal sent to the coil worked as expected. At the moment, we have not identified the origin of the issue. During the maintenance on Tuesday, we will perform a check at WEB to inspect the setup and investigate the problem.

The injections started at ~11:20 UTC and ended at ~18:30 UTC and the output files are available in (/virgoData/NoiseInjections/MagneticInjectionsO4/output):

Far-field sweep magnetic injections

  • CEB--> MagneticSweep_NOISE_MAG_CEB-1453288903.txt
  • NEB --> MagneticSweep_NOISE_MAG_NEB-1453289832.txt

Low frequency lines injection (lines parameters are reported below)

  • NEB--> MagneticLine_NOISE_MAG_NEB-1453290488.txt
  • CEB--> MagneticLine_NOISE_MAG_CEB-1453302290.txt

FreqList = 2.577 3.677 4.677 5.777 6.777 7.477 8.977 10.577 11.977 13.577 15.377 17.577 19.177 22.577 24.577 
DurList = 900 900 900 900 900 900 900 900 600 600 600 600 600 600 600
Amp= 2.0 2.0 2.0 4.0 2.0 2.0 2.0 2.0 2.0 2.0 2.0 2.0 2.0 2.0 2.0

 

**** The interferometer was left in PREPARE SCIENCE, LOW_NOISE_3_ALIGNED with AUTORELOCK_FAILSAFE ****

Images attached to this comment
Lunghini, Fiori, Tringali - 13:34 Tuesday 27 January 2026 (68583) Print this report

This morning we went to WEB to investigate the issue which prevented the injection of magnetic noise on January 24th, 2026. 
We found that the DAC (DAC 1955 Channel 6) for the magnetic injections was connected to an ADC to monitor the output of the DAC in order to investigate the noise problem reported in #68362 and #68372.
We restore the nominal configuration for DAC1955 (see attached Figure 1):
- Channel 6 --> LARGE COIL ANALOG CTRL IN (ENV_NOISE_MAG_WEB channel);
- Channel 7 --> LOUDSPEAKER ANALOG CTRL IN (ENV_NOISE_WEB channel);
We also did a quick set of magnetic and acoustic injection using ENV metatron node and the system worked for both injections type (see attached Figure 2 and Figure 3).

Images attached to this comment
Lunghini, Tringali, Fiori - 1:21 Saturday 21 February 2026 (68729) Print this report

CEB Magnetic Lines Injections started by the operator (Lunghini, #68725) at 20:46 UTC were completed at 00:04 UTC. 
The output file is: 
                            /virgoData/NoiseInjections/MagneticInjectionsO4/output/MagneticLine_NOISE_MAG_CEB-1455655608.txt 

FreqList = 2.577 3.677 4.677 5.777 6.777 7.477 8.977 10.577 11.977 13.577 15.377 17.577 19.177 22.577 24.577 Hz
DurList = 900 900 900 900 900 900 900 900 600 600 600 600 600 600 600 s
Amp = 6.0 6.8 7.4 8.0 8.2 8.2 8.4 8.4 4.0 4.0 4.0 4.0 4.0 4.0 4.0 A

The ITF was left in COMMISSIONING Mode and LOW_NOISE_3_ALIGNED State with AUTORELOCK_FAILSAFE STOP as requested by the experts.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×